Laramie Company produces custom windows to specifications provided by architects. At the end of its accounting period, its account balances indicated the following:
Raw Materials Inventory $76,000
Work in Process Inventory 84,000
Finished Goods Inventory 48,000
Cost of Goods Sold 468,000
Manufacturing Overhead (credit balance) 50,000

Determine the adjusted balances of the accounts if the balance in Manufacturing Overhead is considered immaterial in amount and assigned to Cost of Goods Sold.

Adjusted Balance Raw Material Inventory $ ??
Work in Process Inventory $ ??
Finished Goods Inventory $ ??
Cost of Goods Sold $ ??

Determine the adjusted balances of the accounts if the balance in Manufacturing Overhead is considered material in amount.

Adjusted Balance Raw Material Inventory $ ??
Work in Process Inventory $ ??
Finished Goods Inventory $ ??
Cost of Goods Sold $??
Answer #1

Manufacturing overhead credit balance means overhead is over applied

1) Determine the adjusted balances of the accounts if the balance in Manufacturing Overhead is considered immaterial in amount and assigned to Cost of Goods Sold.

Adjusted Balance Raw Material Inventory $76000
Work in Process Inventory $84000
Finished Goods Inventory $48000
Cost of Goods Sold (468000-50000) = $418000

2) Determine the adjusted balances of the accounts if the balance in Manufacturing Overhead is considered material in amount.

Adjusted Balance Raw Material Inventory $76000
Work in Process Inventory (84000*50000/600000)-84000 = $77000
Finished Goods Inventory (48000*50000/600000)-48000 = $44000
Cost of Goods Sold (468000*50000/600000)-468000 = $429000